What is the method for Conversion of methyl alcohol to aceatdehyde

Grade:12th pass

3 Answers

Shramana Kar
6 years ago
CH3OH + HI & red P = CH3I +KCN = CH3CN + H3O+ (hydrolysis) = CH3COOH + LiAlH4 (in ether medium) = CH3CH2OH + PCC ( pyridinium chloro chromate) = CH3CHO.
SAMANAWAY DAS
6 years ago
always do conversion in the smallest way add PCC with CH3OH you will gate acetaldehyde
CH3OH+PCC=CH3CHO
never ioxidize it will collins reagent or K2Cr2O7 BOTH OF THEM OXDISES ALCOHOL DIRECT TO ORGANIC ACIDS
 
Shramana Kar
6 years ago
CH3OH +PCC = HCHO . It do not forms acetaldehyde. The common reaction is R-CH2OH + Weak oxidizing agent = R-CHO.